# Limb Observations of Global Solar Coronal EUV Wavefronts: the Inclination, Kinematics, Coupling with the Expanding CMEs, and Connection with the CME-driven Shocks

### v1

@@ 3.4 Connection between the EUV Wavefront and the CME-driven Shock
The white-light CME-driven shocks in the corona for all the events have expanded laterally to the side opposite the eruption site on the solar limb. For the unique 2017 September 10 event, the EUV wave is transmitted by the polar coronal holes and is connected with the white-light shock on the opposite side (Hu et al., 2019). For the other five events, the EUV wavefront travels for a distance no larger than 90° (see Column (3) of Table 2), and the white-light shock flank is still traveling after the EUV wavefront has vanished at all altitudes. As noted in Column (7) of Table 2, the white-light shock flank in the later stage is more than 70° away from the final position angle of the EUV wavefront. The position angle of the shock flank is estimated at 2 in the SOHO/LASCO C2 running-difference images, which cannot be geometrically linked by a straight line to the EUV wavefront at 1 . This implies that the EUV wavefront is unlikely connected to the white-light shock flank for the five events, and is no longer the footprint of the shock after the shock has propagated to the opposite side.
@@ 4.1 Inclination of the Primary and Transmitted Wavefronts
The inclination of the primary EUV wavefront is evident on the solar limb in the initial propagation stage for five of the six events (see Figure 1). The wavefront inclination is considered a natural result of the increasing fast magnetosonic speed with the altitude in the low corona, which is applicable to freely propagating EUV waves (e.g., Uchida, 1968; Warmuth et al., 2004; Afanasyev & Uralov, 2011; Kwon & Vourlidas, 2017; Mann & Veronig, 2023). In our events, the inclination of the EUV wavefront occurs in the early stage, in which the CME structure is rapidly expanding and is still coupled with the EUV wavefront. This means that the inclination of the EUV wavefront illustrates the forward inclined loops at the boundary of the expanding CME. If the inclination is caused by the fast magnetosonic speed increasing with the altitude, the inclination angle, defined in this study, should decrease with distance according to models (e.g., Uchida, 1968; Afanasyev & Uralov, 2011; Kwon & Vourlidas, 2017). This suggests that the separation between the time-distance profiles, as displayed in panel (a) of Figure 2, 4–7, should also increase with time. However, except for the 2017 September 10 event, this increase is not seen in the time-distance profiles. For the 2017 September 10 event, the inclination is only observed briefly in the early stage of the propagation, when the CME is still rapidly expanding and is coupled with the EUV wavefront. The inclination angle of the primary wavefront for the five events ranges in – (see Column (5) of Table 1), which reflects the geometry of the forward inclined loops at the boundary of the expanding CME. The geometry may be related to a few factors, such as the initial magnetic configuration of the erupted flux rope and the source region; the eruption direction when the CME has a non-radial motion in the corona (e.g., Möstl et al., 2015; Wang et al., 2015; Hu et al., 2017; Wang et al., 2020), for which the loops on the propagation path may be more inclined toward the solar surface; and the height of the magnetic reconnection that triggers the eruption (i.e., near the eruption site the CME boundary loops expanded by a higher eruption might be more inclined).
@@
4.2 The Speed of the EUV Wavefront and its Comparison with the Shock Lateral Speed
@@ 4.3 Continuous Coupling of the EUV Wavefront and the CME boundary
The interpretation combining the wave and non-wave components of an EUV wave has become more popular, which is supported by the observations of the decoupling of the EUV wavefront and the CME driver (e.g., Patsourakos & Vourlidas, 2009; Cheng et al., 2012; Liberatore et al., 2023). As detailed in Section 3.3, for the 2011 September 22 and 2013 November 7 events, around the time when the EUV wavefront fades out at all altitudes, the CME boundary is observed simultaneously at a similar position angle where the EUV wavefront vanishes (see Column (6) of Table 2). This indicates that the EUV wavefront is coupled with the CME boundary loops for the whole propagation. These two events are different from the other four events in which the EUV wavefront decouples either before or when encountering a local loop system or coronal hole. For the two events on 2011 September 22 and 2013 November 7, in the path of the EUV wavefront propagation there is also a local loop system in the low corona. However, both the EUV wavefront and expanding CME structure traverse the loop system and are coupled until the wavefront disappears. During the lateral expansion of the CME, the boundary loops in the low corona become more inclined and eventually lie on the solar surface, while the loops are stretched straight in the high corona (e.g., above 2 as displayed in Figure 9). As a result, the CME extends laterally to as large as the travel distance of the EUV wavefront. The two events suggest that an EUV wavefront can be continuously driven by the expanding CME and not be free during the whole propagation. For the first time, we observe the continuous coupling of an EUV wavefront with the associated CME throughout the entire wavefront travel. In this condition, the speed of the EUV wavefront is closely related to the lateral expansion of the CME but not to the fast magnetosonic speed, which should be used with caution in coronal seismology.
@@ 4.4 Connection with the EUV Wavefront and the Geometry of the CME-driven Shock
Except for the 2017 September 10 and 2020 November 29 events, after the EUV wavefront has faded out, the CME-driven shock continues propagating to the opposite side of the Sun, which results in a large angular distance () between the shock flank and the EUV wavefront. With the angular distance , geometrically the shock flank at 2 cannot be linked to the EUV wavefront at 1 by a straight line. Considering that the shock flank is convex and bulges to the propagation direction (indicated by the arrow in Figure 9), when the shock flank is on the opposite side of the Sun, it can no longer connect with the EUV wavefront. As suggested by models in which the fast magnetosonic speed increases with the altitude in the low corona (e.g., Uchida, 1968; Afanasyev & Uralov, 2011; Kwon & Vourlidas, 2017; Mann & Veronig, 2023), the inclination angle of a wavefront in the low corona should decrease with the travel distance. At a proper distance the inclination angle is reduced to nearly zero, so the low-corona wavefront crashes onto the solar surface and eventually dissipates. At the beginning a CME-driven shock is usually a closed dome that is connected to the EUV wavefront as the footprint (e.g., Patsourakos & Vourlidas, 2009; Veronig et al., 2010; Li et al., 2012; Cunha-Silva et al., 2018; Mann & Veronig, 2023), and in the late stage the shock has a sphere-like envelope (e.g., Kwon & Vourlidas, 2017; Liu et al., 2017; Zhu et al., 2021). In this study, after the EUV wavefront has vanished and the shock flank has traveled to the side opposite the eruption site, the shock in the low corona dissipates and the closure of the sphere-like shock bubble begins to collapse. This illustrates the final geometry of a large-scale CME-driven shock when it has traveled to the opposite side and has disconnected from the corresponding EUV wavefront. The final angular width of the CME-driven shock is obviously larger than that of the EUV wavefront. This is different from the unique 2017 September 10 event, where the shock is transmitted by polar coronal holes and is connected with the EUV wavefront throughout the propagation (Hu et al., 2019). Additionally, as shown in Figure 9, all the shock flanks are far ahead of the CME boundaries, which means that the laterally expanding shocks are freely propagating on the opposite side and are no longer being driven by their CMEs. A CME-driven shock plays a key role in the longitudinal distribution of energetic particles (e.g., Rouillard et al., 2012; Park et al., 2013; Lario et al., 2016; Kouloumvakos et al., 2016; Zhu et al., 2018). Unveiling the geometry of a global CME-driven shock that has traveled to the other side of the Sun contributes to the understanding of the release and distribution of energetic particles in the heliosphere.
@@ 5 Conclusions
The wavefront speed can be elevated by a local loop system at lower altitudes below 1.15 in three of the six events, which results in distortion of the wavefront in the low corona. In addition to the wavefront inclination, this could make it difficult to estimate the actual speed of an on-disk EUV wavefront from single viewpoint observations.
@@
The average fast magnetosonic Mach numbers are in the range of 1.2–2.1 at altitudes from 1.05 to 1.25 for four of the six events, and their corresponding average speeds are generally in the range of 300–500 km s-1. The Mach number larger than unity indicates that the EUV wavefront is a shock even at the coronal base in the initial stage and is still coupled with the CME.
@@
In four of the six events, after the CME-driven shock has propagated to the opposite side of the Sun, it can be ° away from the position angle where the EUV wavefront vanishes, which suggests that the shock is no longer connected with the EUV wavefront, losing its footprint in the low corona. After a long-distance propagation, the EUV wavefront in the low corona eventually crashes and dissipates on the solar surface.
@@ Appendix A Local Loops in the Low Corona
For the 2013 May 13 event, as shown in Figure 3, the wavefront speed is elevated between position angles 10° and 20°. In the range of the position angle, a possible coronal cavity can be seen in the AIA 211 Å image in Figure 10(b), which may indicate a loop system (like Liu et al., 2012; Gibson, 2018). As shown in Figure 10(a), the PFSS model indeed reveals a loop system in the low corona that is consistent with the cavity. However, the position-angle range of the PFSS-extrapolated loops does not correspond to the range of the 211 Å cavity, where no closed magnetic field lines are obtained below a position angle 15°. This may be due to (1) the magnetogram data on the solar limb used by the PFSS model are not real-time and are taken more than 10 days ago, and (2) the observational uncertainty of magnetogram data in the polar regions impedes the determination of the actual magnetic field structures.

### v2

@@ 3.4 Connection between the EUV Wavefront and the CME-driven Shock
The white-light CME-driven shocks in the corona for all the events have expanded laterally to the side opposite the eruption site on the solar limb. For the unique 2017 September 10 event, the EUV wave is transmitted by the polar coronal holes and is connected with the white-light shock on the opposite side (Hu et al., 2019). For the other five events, the EUV wavefront travels for a distance no larger than 90° (see Column (3) of Table 2), and the white-light shock flank is still traveling after the EUV wavefront has vanished at all altitudes. As noted in Column (7) of Table 2, the white-light shock flank in the later stage is more than 70° away from the final position angle of the EUV wavefront (for the 2020 November 29 event this is true for the southward shock flank as discussed in the previous paragraph). The position angle of the shock flank is estimated at 2 in the SOHO/LASCO C2 running-difference images, which cannot be geometrically linked by a straight line to the EUV wavefront at 1 . This implies that the EUV wavefront is unlikely connected to the white-light shock flank for the five events, and is no longer the footprint of the shock after the shock has propagated to the opposite side.
@@ 4.1 Inclination of the Primary and Transmitted Wavefronts
The inclination of the primary EUV wavefront is evident on the solar limb in the initial propagation stage for five of the six events (see Figure 1). The wavefront inclination is considered a natural result of the increasing fast magnetosonic speed with the altitude in the low corona, which is applicable to freely propagating EUV waves (e.g., Uchida, 1968; Warmuth et al., 2004; Afanasyev & Uralov, 2011; Kwon & Vourlidas, 2017; Mann & Veronig, 2023). In our events, the inclination of the EUV wavefront occurs in the early stage, in which the CME structure is rapidly expanding and is still coupled with the EUV wavefront. This means that the inclination of the EUV wavefront illustrates the forward inclined loops at the boundary of the expanding CME. If the inclination is caused by the fast magnetosonic speed increasing with the altitude, the inclination angle, defined in this study, should decrease with distance according to models (e.g., Uchida, 1968; Afanasyev & Uralov, 2011; Kwon & Vourlidas, 2017). This requires that the separation between the time-distance profiles, as displayed in panel (a) of Figure 2, 4–7, should also increase with time. However, except for the 2017 September 10 event, this increase is not seen in the time-distance profiles. For the 2017 September 10 event, the inclination is only observed briefly in the early stage of the propagation, when the CME is still rapidly expanding and is coupled with the EUV wavefront. The inclination angle of the primary wavefront for the five events ranges in – (see Column (5) of Table 1), which reflects the geometry of the forward inclined loops at the boundary of the expanding CME. The geometry may be related to a few factors, such as the initial magnetic configuration of the erupted flux rope and the source region; the eruption direction when the CME has a non-radial motion in the corona (e.g., Möstl et al., 2015; Wang et al., 2015; Hu et al., 2017; Wang et al., 2020), for which the loops on the propagation path may be more inclined toward the solar surface; and the height of the magnetic reconnection that triggers the eruption (i.e., near the eruption site the CME boundary loops expanded by a higher eruption might be more inclined).
@@
4.2 The Speed of the EUV Wavefront and its Comparison with the Shock Lateral Expansion Speed
@@ 4.3 Continuous Coupling of the EUV Wavefront and the CME boundary
The interpretation combining the wave and non-wave components of an EUV wave has become more popular, which is supported by the observations of the decoupling of the EUV wavefront and the CME driver (e.g., Patsourakos & Vourlidas, 2009; Cheng et al., 2012; Liberatore et al., 2023). As detailed in Section 3.3, for the 2011 September 22 and 2013 November 7 events, around the time when the EUV wavefront fades out at all altitudes, the CME boundary is observed simultaneously at a similar position angle where the EUV wavefront vanishes (see Column (6) of Table 2). This indicates that the EUV wavefront is coupled with the CME boundary loops for the whole propagation. These two events are different from the other four events in which the EUV wavefront decouples either before or when encountering a local loop system or coronal hole. For the two events on 2011 September 22 and 2013 November 7, in the path of the EUV wavefront propagation there is also a local loop system in the low corona. However, both the EUV wavefront and expanding CME structure traverse the loop system and are coupled until the wavefront disappears. During the lateral expansion of the CME, the boundary loops in the low corona become more inclined and eventually lie on the solar surface, while the loops are stretched straight in the high corona (e.g., above 2 as displayed in Figure 9). As a result, the CME extends laterally to as large as the travel distance of the EUV wavefront. The two events suggest that an EUV wavefront can be continuously driven by the expanding CME and not be free during the whole propagation. For the first time, we observe the continuous coupling of an EUV wavefront with the associated CME throughout the entire wavefront travel. In this condition, the speed of the EUV wavefront is closely related to the lateral expansion of the CME but not to the local fast magnetosonic speed, which should be used with caution in coronal seismology.
@@ 4.4 Connection with the EUV Wavefront and the Geometry of the CME-driven Shock
Except for the 2017 September 10 and 2020 November 29 events, after the EUV wavefront has faded out, the CME-driven shock continues propagating to the opposite side of the Sun, which results in a large angular distance () between the shock flank and the EUV wavefront (actually for the 2020 November 29 event the angular distance for the southward shock flank is also ). With the angular distance , geometrically the shock flank at 2 cannot be linked to the EUV wavefront at 1 by a straight line. Considering that the shock flank is convex and bulges to the propagation direction (indicated by the arrow in Figure 9), when the shock flank is on the opposite side of the Sun, it can no longer connect with the EUV wavefront. As suggested by models in which the fast magnetosonic speed increases with the altitude in the low corona (e.g., Uchida, 1968; Afanasyev & Uralov, 2011; Kwon & Vourlidas, 2017; Mann & Veronig, 2023), the inclination angle of a free wavefront in the low corona should decrease with the travel distance. At a proper distance the inclination angle is reduced to nearly zero, so the low-corona wavefront crashes onto the solar surface and eventually dissipates. At the beginning a CME-driven shock is usually a closed dome that is connected to the EUV wavefront as the footprint (e.g., Patsourakos & Vourlidas, 2009; Veronig et al., 2010; Li et al., 2012; Cunha-Silva et al., 2018; Mann & Veronig, 2023), and in the late stage the shock has a sphere-like envelope (e.g., Kwon & Vourlidas, 2017; Liu et al., 2017; Zhu et al., 2021). In this study, after the EUV wavefront has vanished and the shock flank has traveled to the side opposite the eruption site, the shock in the low corona dissipates and the closure of the sphere-like shock bubble begins to collapse. This illustrates the final geometry of a large-scale CME-driven shock when it has traveled to the opposite side and has disconnected from the corresponding EUV wavefront. The final angular width of the CME-driven shock is obviously larger than that of the EUV wavefront. This is different from the unique 2017 September 10 event, where the shock is transmitted by polar coronal holes and is connected with the EUV wavefront throughout the propagation (Hu et al., 2019). Additionally, as shown in Figure 9, all the shock flanks are far ahead of the CME boundaries, which means that the laterally expanding shocks are freely propagating on the opposite side and are no longer being driven by their CMEs. A CME-driven shock plays a key role in the longitudinal distribution of energetic particles (e.g., Rouillard et al., 2012; Park et al., 2013; Lario et al., 2016; Kouloumvakos et al., 2016; Zhu et al., 2018). Unveiling the geometry of a global CME-driven shock that has traveled to the other side of the Sun contributes to the understanding of the release and distribution of energetic particles in the heliosphere.
@@ 5 Conclusions
The wavefront speed is elevated by a local loop system only at lower altitudes below 1.15 in three of the six events, which results in distortion of the wavefront in the low corona. In addition to the wavefront inclination, this could make it difficult to estimate the actual speed of an on-disk EUV wavefront from single viewpoint observations.
@@
The average fast magnetosonic Mach numbers are in the range of 1.2–2.1 at altitudes from 1.05 to 1.25 for four of the six events, and their corresponding average speeds are generally in the range of 300–500 km s-1. The Mach number larger than unity indicates that the EUV wavefront is a shock even at the coronal base in the initial stage when the EUV wavefront is still coupled with the CME.
@@
In four of the six events, after the CME-driven shock has propagated to the opposite side of the Sun, it can be ° away from the position angle where the EUV wavefront vanishes, which suggests that the shock is no longer connected with the EUV wavefront, losing its footprint in the low corona. After a long-distance propagation, the CME-driven shock in the low corona eventually crashes and dissipates on the solar surface.
@@ Appendix A Local Loops in the Low Corona
For the 2013 May 13 event, as shown in Figure 3, the wavefront speed is elevated between position angles and . In the range of the position angle, a possible coronal cavity can be seen in the AIA 211 Å image in Figure 10(b), which may indicate a loop system (like Liu et al., 2012; Gibson, 2018). As shown in Figure 10(a), the PFSS model indeed reveals a loop system in the low corona that is consistent with the cavity. However, the position-angle range of the PFSS-extrapolated loops does not correspond to the range of the 211 Å cavity, where no closed magnetic field lines are obtained below a position angle 15°. This may be due to (1) the magnetogram data on the solar limb used by the PFSS model are not real-time and are taken more than 10 days ago, and (2) the observational uncertainty of magnetogram data in the polar regions impedes the determination of the actual magnetic field structures.